Calculate the pKa of a 0.062 M HOCl solution with a pH of 5.84. 

Provide answer to two decimal places and without units.



1
Expert's answer
2021-03-17T06:46:22-0400

HOCl = H+ + OCl-

Ka = [H+][OCl-]/[HOCl] = [H+]2/(c(HOCl) - [H+]) = 10-2*5.84/(0.062 - 10-5.84) = 3.37*10-11

pKa = -log(Ka) = 10.47